# Lie algebras generated by reflections in types BCD

### v1

@@ 2.5. Restriction of Type B simple modules
see [16, Theorem 4.1] or [15, Theorem 6.6]. In terms of weight spaces, (2.5.1) takes the form:
@@ 3.1. Type D Weyl groups
Recall the linear characters $\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}$ on $\mathcal{B}_{n}$ defined in (2.2.1). Set $\mathcal{D}_{n}=\ker(\varepsilon^{\prime})$ and $N_{n}^{\prime}=N_{n}\cap\mathcal{D}_{n}$. Let $\widetilde{s}_{n}=t_{n}s_{n-1}t_{n}=t_{n-1}t_{n}s_{n-1}\in\mathcal{D}_{n}$. Then for $n\geq 4$, the pair $(\mathcal{D}_{n},\left\{s_{1},\ldots,s_{n-1},\widetilde{s}_{n}\right\})$ is a Coxeter system of type $D_{n}$, and $\varepsilon^{\prime\prime}$ restricts to the sign character of $\mathcal{D}_{n}$. One has $\mathcal{D}_{n}=N_{n}^{\prime}\rtimes H_{n}$, and each element of $\mathcal{D}_{n}$ is uniquely of the form $t_{1}^{i_{1}}\cdots t_{n}^{i_{n}}\sigma$ with $\sigma\in H_{n}$, $i_{j}\in\left\{0,1\right\}$, and $i_{1}+\cdots+i_{n}$ even. For $11\leq i\leq n-$ 1, let $u_{i}=t_{1}t_{i+1}$. Then each element of $\mathcal{D}_{n}$ is also uniquely of the form $u_{1}^{i_{1}}\cdots u_{n-1}^{i_{n-1}}\sigma$ with $\sigma\in H_{n}$ and $i_{j}\in\left\{0,1\right\}$ for each $j$. One has $N_{n}^{\prime}\cong\mathbb{Z}_{2}^{n-1}$; via this identification, $u_{i}$ is a multiplicative generator for the $i$-th factor in $\mathbb{Z}_{2}^{n-1}$. Identifying $H_{n}$ with $\mathfrak{S}_{n}$, the set of reflections in $\mathcal{D}_{n}$ is
@@ 3.2. Simple modules for Type D Weyl groups
be the set of all unordered pairs of partitions whose parts together sum to $n$. The simple $\mathbb{C}\mathcal{D}_{n}$-modules are now obtained as follows; cf. [5, Proposition 5.1].
@@
By [8, Corollary 5.6.4], the field $\mathbb{Q}$ is a splitting field for $\mathcal{D}_{n}$. Then all irreducible characters of $\mathcal{D}_{n}$ are real valued, and hence all simple $\mathbb{C}\mathcal{D}_{n}$-modules are self-dual.
@@
For a given value of $n$, this labeling of $S^{\left\{\lambda,\pm\right\}}$ may be the opposite of that corresponding to the characters $\chi_{(\lambda,\pm)}$ defined in [8, §5.6] or [7, §3]. For $T\in\mathbb{T}(\lambda,\lambda)$, set $S_{T}^{\left\{\lambda,\pm\right\}}=\operatorname{span}\{c_{T}\pm c_{T^{\natural}}\}$. Define subsets $\mathbb{T}(\lambda,\lambda)_{\pm 1}$ of $\mathbb{T}(\lambda,\lambda)$ by $\mathbb{T}(\lambda,\lambda)_{\pm 1}=\{T\in\mathbb{T}(\lambda,\lambda):\rho_{T}(1)=\pm 1\}$. Then $\mathbb{T}(\lambda,\lambda)$ is the disjoint union of the sets $\mathbb{T}(\lambda,\lambda)_{+1}$ and $\mathbb{T}(\lambda,\lambda)_{-1}$, and the sets
@@
cf. [7, Lemma 3.5] (where the labeling of characters may be opposite to our labeling).
@@ 3.3. Weight space decompositions of Type D simple modules
where $\widetilde{\rho}_{i}\coloneq\rho_{1}\rho_{i+1}$ is the eigenvalue by which $u_{i}=t_{1}t_{i+1}$ acts on $V_{\alpha}$. One immediately checks that if $\alpha,\beta\in\mathcal{W}(n)$ with $\alpha=(\rho,\chi)$, then $\widetilde{\alpha}=\widetilde{\beta}$ if and only if $\beta=\alpha$ or $\beta=(-\rho,\chi)$. Combined with (2.3.3), this implies for $(\lambda,\mu)\in\mathcal{BP}(n)$ that the sets $\mathcal{W}(\lambda,\mu)$ and $\mathcal{W}(\mu,\lambda)$ have the same restrictions to $GZ(\mathcal{D}_{n})$. Then we can unambiguously define $\mathcal{W}\{\lambda,\mu\}\coloneq\left\{\widetilde{\alpha}:\alpha\in\mathcal{W}(\lambda,\mu)\right\}$. One further deduces that the union $\mathcal{W}\{n\}\coloneq\bigcup_{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}}\mathcal{W}\{\lambda,\mu\}$ is disjoint. One sees for $\lambda\neq\mu$ that the weights in the set $\mathcal{W}(\lambda,\mu)$ restrict to distinct elements in $\mathcal{W}\{\lambda,\mu\}$, while for $\lambda=\mu$, the weights in $\mathcal{W}(\lambda,\lambda)$ occur in pairs (±$(\pm\rho,\chi)$ that have common restrictions to $GZ(\mathcal{D}_{n})$.
@@ 3.4. Orthogonal forms
    Suppose $i$ and $i+1$ are not in the same tableau in $T$. Let $S=s_{i}\cdot T$. Then with respect to the basis $\left\{c_{T}\pm c_{T^{\natural}},c_{S}\pm c_{S^{\natural}}\right\}$, $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}\oplus S_{S}^{\left\{\lambda,\pm\right\}}$ via the matrix $[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}]$.
@@
    If $i$ and $i+1$ are in the same column of the same tableau in $T$, implying that $r_{i}=-1$, then $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}$ as multiplication by $-1$.
@@
    If $i$ and $i+1$ are in the same row of the same tableau in $T$, implying that $r_{i}=+1$, then $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}$ as multiplication by $+1$.
@@
In the case $i=1$ of the lemma, if $1$ and $2$ are not in the same tableau in $T$ and if $S=s_{1}\cdot T$, then $S\notin\mathbb{T}(\lambda,\lambda)_{+1}$ but rather $S^{\natural}\in\mathbb{T}(\lambda,\lambda)_{+1}$. Then $c_{S^{\natural}}-c_{S}=-(c_{S}-c_{S^{\natural}})$ is the preferred basis vector for $S_{S^{\natural}}^{\left\{\lambda,-\right\}}$ identified in (3.2.2), and with respect to the basis $\left\{c_{T}-c_{T^{\natural}},c_{S^{\natural}}-c_{S}\right\}$, the element $s_{1}$ acts on $S_{T}^{\left\{\lambda,-\right\}}\oplus S_{S^{\natural}}^{\left\{\lambda,-\right\}}$ via the matrix $[\begin{smallmatrix}\hfill 0&-1\\ -1&\hfill 0\end{smallmatrix}]$.
@@ 3.5. Restriction of Type D simple modules
Lemma 3.5.1 ([17, Theorem and Corollary 2.1]).
@@ 4. Lie algebras generated by reflections
We begin in Section 4.1 by recording some straightforward generalizations of results originally stated in the context of symmetric groups by Marin; our primary sources are [11, III.3] and [13, §2]. In Section 4.2 we introduce the Lie algebras $\mathfrak{b}_{n}$ and $\mathfrak{d}_{n}$ that are the main objects of study in this paper, and then in Section 4.3 we begin to describe the images of $\mathfrak{b}_{n}$ and $\mathfrak{d}_{n}$ in the various factors on the right-hand sides of the Artin–Wedderburn isomorphisms (2.2.4) and (3.2.3).
@@ 4.1. Generalities
Let $G$ be a finite group and let $\emptyset\neq S\subseteq G$. Write $\operatorname{Lie}(\mathbb{C}G)$ for the group algebra $\mathbb{C}G$ considered as a Lie algebra via the commutator bracket $[x,y]=xy-yx$. Let $\mathfrak{g}$ be the Lie subalgebra of $\operatorname{Lie}(\mathbb{C}G)$ generated by the elements of the set $S$, let $\mathfrak{g}^{\prime}=[\mathfrak{g},\mathfrak{g}]$ be the derived subalgebra of $\mathfrak{g}$, and let $Z(\mathfrak{g})$ be the center of $\mathfrak{g}$. Throughout this section we make the following assumptions (though in practice (G3) can be weakened; see Remark 4.1.7):
@@
Let $V_{1},V_{2},\ldots,V_{m}$ be a complete set of pairwise non-isomorphic simple $\mathbb{C}G$-modules. Then $V:=\bigoplus_{i=1}^{m}V$ is a faithful finite-dimensional $\mathbb{C}G$-module. Since $\mathfrak{g}$ contains the set $S$ of associative algebra generators for $\mathbb{C}G$, it follows that each $V_{i}$ restricts to a simple $\mathfrak{g}$-module, and hence $V$ restricts to a faithful, finite-dimensional, semisimple $\mathfrak{g}$-module. Then $\mathfrak{g}$ is reductive, $\mathfrak{g}^{\prime}$ is semisimple, and $\mathfrak{g}=\mathfrak{g}^{\prime}\oplus Z(\mathfrak{g})$ by Proposition 5 of [1, Chapter I, §6, no. 4]. ∎
@@
For $1\leq i\leq t$, one has $T_{i}\in Z(\mathfrak{g})$ by Lemma 4.1.1. Conversely, if $z\in Z(\mathfrak{g})$, then $z\in Z(\mathbb{C}G)$ by the lemma, and hence $z$ is a linear combination of class sums in $\mathbb{C}G$. Given a conjugacy class $C$ in $G$, let $\Psi_{C}:\mathbb{C}G\to\mathbb{C}$ be the (linear extension to $\mathbb{C}G$ of the) class function that evaluates to $1$ on the elements of $C$ and that evaluates to 00 on the other elements of $G$. If $0\neq z\in Z(\mathbb{C}G)$, then $\Psi_{C}(z)\neq 0$ for some conjugacy class $C$.
Given a sequence $E=(s_{1},\ldots,s_{r})$ of elements of the set $S$, set $\ell(E)=r$, set $s_{E}=s_{1}$ if $r=1$, and set $s_{E}=[s_{1},s_{E^{\prime}}]$ if $r>1$ and $E^{\prime}=(s_{2},\ldots,s_{r})$. Then $\mathfrak{g}$ is spanned by the elements $s_{E}$ as $E$ ranges over all finite sequences of elements in $S$, and $\mathfrak{g}^{\prime}$ is spanned by all $s_{E}$ with $\ell(E)>1$. If $\ell(E)>1$ and $s=s_{1}$, then $s_{E}=[s,s_{E^{\prime}}]=ss_{E^{\prime}}-s_{E^{\prime}}s$ can be written as a linear combination of terms of the form $sg-gs$ with $g\in G$. Since the elements of $S$ are of order $2$, one has $gs=s(sg)s^{-1}$. Then $gs$ and $sg$ are conjugate in $G$, and hence $\Psi_{C}([s,g])=\Psi_{C}(sg)-\Psi_{C}(gs)=0$. It follows for each conjugacy class $C$ of $G$ that $\Psi_{C}$ vanishes on $\mathfrak{g}^{\prime}$, and $\Psi_{C}$ is nonzero on $\mathfrak{g}$ only if $C=C_{i}$ for some $1\leq i\leq t$. Thus if $0\neq z\in Z(\mathfrak{g})$, then $z$ must be a linear combination of $T_{1},\ldots,T_{t}$. ∎
@@
If $V$ is a $\mathbb{C}G$-module, then the dual space $V^{*}=\operatorname{Hom}(V,\mathbb{C})$ is a $\mathbb{C}G$-module with action defined for $g\in G$, $\phi\in V^{*}$, and $v\in V$ by $(g.\phi)(v)=\phi(g^{-1}.v)$. On the other hand, $[\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V)]^{*}$ is a $\mathfrak{g}$-module via the contragredient action, defined for $x\in\mathfrak{g}$, $\phi\in V^{*}$, and $v\in V$ by $(x.\phi)(v)=-\phi(x.v)$; we denote $V^{*}$ with this $\mathfrak{g}$-module structure by $V^{*,\operatorname{Lie}}$. In general, the $\mathfrak{g}$-modules $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V^{*})$ and $V^{*,\operatorname{Lie}}$ need not be isomorphic.
@@
Let $\kappa:G\to\left\{\pm 1\right\}$ be a linear character that evaluates to $-1$ on each element of $S$. Then with notation and assumptions as above, $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V^{*}\otimes\kappa)\cong V^{*,\operatorname{Lie}}$ as $\mathfrak{g}$-modules. If $V\cong V^{*}$ as $\mathbb{C}G$-modules, then $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V\otimes\kappa)\cong V^{*,\operatorname{Lie}}$ as $\mathfrak{g}$-modules.

### v2

@@ 2.5. Restriction of Type B simple modules
see [Okada:1990, Theorem 4.1] or [Mishra:2016, Theorem 6.6]. In terms of weight spaces, (2.5.1) takes the form:
@@ 3.1. Type D Weyl groups
Recall the linear characters $\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}$ on $\mathcal{B}_{n}$ defined in (2.2.1). Set $\mathcal{D}_{n}=\ker(\varepsilon^{\prime})$ and $N_{n}^{\prime}=N_{n}\cap\mathcal{D}_{n}$. Let $\widetilde{s}_{n}=t_{n}s_{n-1}t_{n}=t_{n-1}t_{n}s_{n-1}\in\mathcal{D}_{n}$. Then for $n\geq 4$, the pair $(\mathcal{D}_{n},\left\{s_{1},\ldots,s_{n-1},\widetilde{s}_{n}\right\})$ is a Coxeter system of type $D_{n}$, and $\varepsilon^{\prime\prime}$ restricts to the sign character of $\mathcal{D}_{n}$. One has $\mathcal{D}_{n}=N_{n}^{\prime}\rtimes H_{n}$, and each element of $\mathcal{D}_{n}$ is uniquely of the form $t_{1}^{i_{1}}\cdots t_{n}^{i_{n}}\sigma$ with $\sigma\in H_{n}$, $i_{j}\in\left\{0,1\right\}$, and $i_{1}+\cdots+i_{n}$ even. For $1\leq i\leq n-1$, let $u_{i}=t_{1}t_{i+1}$. Then each element of $\mathcal{D}_{n}$ is also uniquely of the form $u_{1}^{i_{1}}\cdots u_{n-1}^{i_{n-1}}\sigma$ with $\sigma\in H_{n}$ and $i_{j}\in\left\{0,1\right\}$ for each $j$. One has $N_{n}^{\prime}\cong\mathbb{Z}_{2}^{n-1}$; via this identification, $u_{i}$ is a multiplicative generator for the $i$-th factor in $\mathbb{Z}_{2}^{n-1}$. Identifying $H_{n}$ with $\mathfrak{S}_{n}$, the set of reflections in $\mathcal{D}_{n}$ is
@@ 3.2. Simple modules for Type D Weyl groups
be the set of all unordered pairs of partitions whose parts together sum to $n$. The simple $\mathbb{C}\mathcal{D}_{n}$-modules are now obtained as follows; cf. [Fulton:1991, Proposition 5.1].
@@
By [Geck:2000, Corollary 5.6.4], the field $\mathbb{Q}$ is a splitting field for $\mathcal{D}_{n}$. Then all irreducible characters of $\mathcal{D}_{n}$ are real valued, and hence all simple $\mathbb{C}\mathcal{D}_{n}$-modules are self-dual.
@@
For a given value of $n$, this labeling of $S^{\left\{\lambda,\pm\right\}}$ may be the opposite of that corresponding to the characters $\chi_{(\lambda,\pm)}$ defined in [Geck:2000, §5.6] or [Geck:2015, §3]. For $T\in\mathbb{T}(\lambda,\lambda)$, set $S_{T}^{\left\{\lambda,\pm\right\}}=\operatorname{span}\{c_{T}\pm c_{T^{\natural}}\}$. Define subsets $\mathbb{T}(\lambda,\lambda)_{\pm 1}$ of $\mathbb{T}(\lambda,\lambda)$ by $\mathbb{T}(\lambda,\lambda)_{\pm 1}=\{T\in\mathbb{T}(\lambda,\lambda):\rho_{T}(1)=\pm 1\}$. Then $\mathbb{T}(\lambda,\lambda)$ is the disjoint union of the sets $\mathbb{T}(\lambda,\lambda)_{+1}$ and $\mathbb{T}(\lambda,\lambda)_{-1}$, and the sets
@@
cf. [Geck:2015, Lemma 3.5] (where the labeling of characters may be opposite to our labeling).
@@ 3.3. Weight space decompositions of Type D simple modules
where $\widetilde{\rho}_{i}\coloneq\rho_{1}\rho_{i+1}$ is the eigenvalue by which $u_{i}=t_{1}t_{i+1}$ acts on $V_{\alpha}$. One immediately checks that if $\alpha,\beta\in\mathcal{W}(n)$ with $\alpha=(\rho,\chi)$, then $\widetilde{\alpha}=\widetilde{\beta}$ if and only if $\beta=\alpha$ or $\beta=(-\rho,\chi)$. Combined with (2.3.3), this implies for $(\lambda,\mu)\in\mathcal{BP}(n)$ that the sets $\mathcal{W}(\lambda,\mu)$ and $\mathcal{W}(\mu,\lambda)$ have the same restrictions to $GZ(\mathcal{D}_{n})$. Then we can unambiguously define $\mathcal{W}\{\lambda,\mu\}\coloneq\left\{\widetilde{\alpha}:\alpha\in\mathcal{W}(\lambda,\mu)\right\}$. One further deduces that the union $\mathcal{W}\{n\}\coloneq\bigcup_{\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}}\mathcal{W}\{\lambda,\mu\}$ is disjoint. One sees for $\lambda\neq\mu$ that the weights in the set $\mathcal{W}(\lambda,\mu)$ restrict to distinct elements in $\mathcal{W}\{\lambda,\mu\}$, while for $\lambda=\mu$, the weights in $\mathcal{W}(\lambda,\lambda)$ occur in pairs $(\pm\rho,\chi)$ that have common restrictions to $GZ(\mathcal{D}_{n})$.
@@ 3.4. Orthogonal forms
    Suppose $i$ and $i+1$ are not in the same tableau in $T$. Let $S=s_{i}\cdot T$. Then with respect to the basis $\left\{c_{T}\pm c_{T^{\natural}},c_{S}\pm c_{S^{\natural}}\right\}$, $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}\oplus S_{S}^{\left\{\lambda,\pm\right\}}$ via the matrix $\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}]$.
@@
    If $i$ and $i+1$ are in the same column of the same tableau in $T$, implying that $r_{i}=-1$, then $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}$ as multiplication by 1-1.
@@
    If $i$ and $i+1$ are in the same row of the same tableau in $T$, implying that $r_{i}=+1$, then $s_{i}$ acts on $S_{T}^{\left\{\lambda,\pm\right\}}$ as multiplication by +1+1.
@@
In the case $i=1$ of the lemma, if $1$ and $2$ are not in the same tableau in $T$ and if $S=s_{1}\cdot T$, then $S\notin\mathbb{T}(\lambda,\lambda)_{+1}$ but rather $S^{\natural}\in\mathbb{T}(\lambda,\lambda)_{+1}$. Then $c_{S^{\natural}}-c_{S}=-(c_{S}-c_{S^{\natural}})$ is the preferred basis vector for $S_{S^{\natural}}^{\left\{\lambda,-\right\}}$ identified in (3.2.2), and with respect to the basis $\left\{c_{T}-c_{T^{\natural}},c_{S^{\natural}}-c_{S}\right\}$, the element $s_{1}$ acts on $S_{T}^{\left\{\lambda,-\right\}}\oplus S_{S^{\natural}}^{\left\{\lambda,-\right\}}$ via the matrix $\begin{smallmatrix}\hfill 0&-1\\ -1&\hfill 0\end{smallmatrix}]$.
@@ 3.5. Restriction of Type D simple modules
Lemma 3.5.1 ([Tokuyama:1984, Theorem and Corollary 2.1]).
@@ 4. Lie algebras generated by reflections
We begin in Section 4.1 by recording some straightforward generalizations of results originally stated in the context of symmetric groups by Marin; our primary sources are [Marin:2001, III.3] and [Marin:2007, §2]. In Section 4.2 we introduce the Lie algebras $\mathfrak{b}_{n}$ and $\mathfrak{d}_{n}$ that are the main objects of study in this paper, and then in Section 4.3 we begin to describe the images of $\mathfrak{b}_{n}$ and $\mathfrak{d}_{n}$ in the various factors on the right-hand sides of the Artin–Wedderburn isomorphisms (2.2.4) and (3.2.3).
@@ 4.1. Generalities
Let $G$ be a finite group and let $\emptyset\neq S\subseteq G$. Write $\operatorname{Lie}(\mathbb{C}G)$ for the group algebra $\mathbb{C}G$ considered as a Lie algebra via the commutator bracket [x,y]=xyyx[x,y]=xy-yx. Let $\mathfrak{g}$ be the Lie subalgebra of $\operatorname{Lie}(\mathbb{C}G)$ generated by the elements of the set $S$, let $\mathfrak{g}^{\prime}=[\mathfrak{g},\mathfrak{g}]$ be the derived subalgebra of $\mathfrak{g}$, and let $Z(\mathfrak{g})$ be the center of $\mathfrak{g}$. Throughout this section we make the following assumptions (though in practice (G3) can be weakened; see Remark 4.1.7):
@@
Let $V_{1},V_{2},\ldots,V_{m}$ be a complete set of pairwise non-isomorphic simple $\mathbb{C}G$-modules. Then $V:=\bigoplus_{i=1}^{m}V$ is a faithful finite-dimensional $\mathbb{C}G$-module. Since $\mathfrak{g}$ contains the set $S$ of associative algebra generators for $\mathbb{C}G$, it follows that each $V_{i}$ restricts to a simple $\mathfrak{g}$-module, and hence $V$ restricts to a faithful, finite-dimensional, semisimple $\mathfrak{g}$-module. Then $\mathfrak{g}$ is reductive, $\mathfrak{g}^{\prime}$ is semisimple, and $\mathfrak{g}=\mathfrak{g}^{\prime}\oplus Z(\mathfrak{g})$ by Proposition 5 of [Bourbaki:1998, Chapter I, §6, no. 4]. ∎
@@
For $1\leq i\leq t$, one has $T_{i}\in Z(\mathfrak{g})$ by Lemma 4.1.1. Conversely, if $z\in Z(\mathfrak{g})$, then $z\in Z(\mathbb{C}G)$ by the lemma, and hence $z$ is a linear combination of class sums in $\mathbb{C}G$. Given a conjugacy class $C$ in $G$, let $\Psi_{C}:\mathbb{C}G\to\mathbb{C}$ be the (linear extension to $\mathbb{C}G$ of the) class function that evaluates to $1$ on the elements of $C$ and that evaluates to 0 on the other elements of $G$. If $0\neq z\in Z(\mathbb{C}G)$, then $\Psi_{C}(z)\neq 0$ for some conjugacy class $C$.
Given a sequence $E=(s_{1},\ldots,s_{r})$ of elements of the set $S$, set $\ell(E)=r$, set $s_{E}=s_{1}$ if $r=1$, and set $s_{E}=[s_{1},s_{E^{\prime}}]$ if $r>1$ and $E^{\prime}=(s_{2},\ldots,s_{r})$. Then $\mathfrak{g}$ is spanned by the elements $s_{E}$ as $E$ ranges over all finite sequences of elements in $S$, and $\mathfrak{g}^{\prime}$ is spanned by all $s_{E}$ with $\ell(E)>1$. If $\ell(E)>1$ and $s=s_{1}$, then $s_{E}=[s,s_{E^{\prime}}]=ss_{E^{\prime}}-s_{E^{\prime}}s$ can be written as a linear combination of terms of the form sggssg-gs with $g\in G$. Since the elements of $S$ are of order $2$, one has $gs=s(sg)s^{-1}$. Then $gs$ and $sg$ are conjugate in $G$, and hence $\Psi_{C}([s,g])=\Psi_{C}(sg)-\Psi_{C}(gs)=0$. It follows for each conjugacy class $C$ of $G$ that $\Psi_{C}$ vanishes on $\mathfrak{g}^{\prime}$, and $\Psi_{C}$ is nonzero on $\mathfrak{g}$ only if $C=C_{i}$ for some $1\leq i\leq t$. Thus if $0\neq z\in Z(\mathfrak{g})$, then $z$ must be a linear combination of $T_{1},\ldots,T_{t}$. ∎
@@
If $V$ is a $\mathbb{C}G$-module, then the dual space $V^{*}=\operatorname{Hom}(V,\mathbb{C})$ is a $\mathbb{C}G$-module with action defined for $g\in G$, $\phi\in V^{*}$, and $v\in V$ by $(g.\phi)(v)=\phi(g^{-1}.v)$. On the other hand, $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V)]^{*}$ is a $\mathfrak{g}$-module via the contragredient action, defined for $x\in\mathfrak{g}$, $\phi\in V^{*}$, and $v\in V$ by $(x.\phi)(v)=-\phi(x.v)$; we denote $V^{*}$ with this $\mathfrak{g}$-module structure by $V^{*,\operatorname{Lie}}$. In general, the $\mathfrak{g}$-modules $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V^{*})$ and $V^{*,\operatorname{Lie}}$ need not be isomorphic.
@@
Let $\kappa:G\to\left\{\pm 1\right\}$ be a linear character that evaluates to 1-1 on each element of $S$. Then with notation and assumptions as above, $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V^{*}\otimes\kappa)\cong V^{*,\operatorname{Lie}}$ as $\mathfrak{g}$-modules. If $V\cong V^{*}$ as $\mathbb{C}G$-modules, then $\operatorname{Res}^{\mathbb{C}G}_{\mathfrak{g}}(V\otimes\kappa)\cong V^{*,\operatorname{Lie}}$ as $\mathfrak{g}$-modules.